# Line bundles and exact sequences for the ideal class group and the Picard group

### v1

@@ 3. Ideal class group of an extension vs Picard group
Let $A\subseteq B$ be an extension of rings. The group $\mathscr{G}(A,B)$ modulo $H=\{Ax:x\in B^{\ast}\}$, the subgroup of principal invertible ideals of this extension, is called the ideal class group of this extension and is denoted by $\mathfrak{C}(A,B)$. Therefore we have the following short exact sequence of Abelian groups:
@@
$\mathbf{(ii)}$ $L$ is a finitely generated projective $A$-module of rank 1 with $LB=B$.
@@
Note that Lemmas 3.3 and 3.5 also improve [8, Lemmas 2.2, 2.3].
@@
For any ring $A$ we have the following exact sequence of Abelian groups: ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{C}(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Pic}(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Pic}(T(A)).}$
@@ 4. The geometry of line bundles
Let $A$ be a ring and $L\in\operatorname{Pic}(A)$. If $f:L\rightarrow A$ is a morphism of $A$-modules, then $f(x)y=f(y)x$ for all $x,y\in L$. In particular, the $A$-algebra $\bigoplus\limits_{n\in\mathbb{Z}}L^{\otimes n}$ is a commutative ring.
@@
It follows from Theorems 3.8 and 4.2. As an alternative proof, in the proof of Theorem 4.2, we observed that for any ring $A$ then there exists an extension ring $A_{1}$ of $A$ such that the induced map $\operatorname{Pic}(A)\rightarrow\operatorname{Pic}(A_{1})$ is zero. Then by Theorem 3.8, we have $\operatorname{Pic}(A)\simeq\mathfrak{C}(A,A_{1})$. ∎
@@
By Theorem 4.2, there exists an extension ring $B$ of $A$ such that $\operatorname{Pic}(B)=0$. Then by Theorem 3.8, there exists an invertible ideal $L^{\prime}$ of the extension $A\subseteq B$ such that $L\simeq L^{\prime}$ as $A$-modules. Then by Lemma 4.9, $L=L_{k}$ for some $k$. ∎
@@ 5. The ideal class group behavior with a tower of extensions
where the rows are exact (see Theorem 3.8) and the second and third vertical maps are the canonical isomorphisms. The above diagram is also commutative, since the commutativity of the left-hand side follows from Corollary 3.6, and the right-hand side is clear. Then by the five lemma, the first vertical map is an isomorphism. ∎
@@
where the rows are exact and the second vertical map is an isomorphism. Hence, the canonical map $\operatorname{\mathfrak{C}}(A)\rightarrow\operatorname{\mathfrak{C}}(A_{\mathrm{red}})$ is injective. However, unlike the Picard group, the canonical injection $\operatorname{\mathfrak{C}}(A)\rightarrow\operatorname{\mathfrak{C}}(A_{\mathrm{red}})$ is not necessarily an isomorphism in general (see Example 5.7).
@@
where the rows are exact (see Corollary 3.10) and the second vertical map is an isomorphism. If the third vertical map is injective, then by the five lemma, the canonical group map $\operatorname{\mathfrak{C}}(A)\rightarrow\operatorname{\mathfrak{C}}(A_{\mathrm{red}})$ will be an isomorphism. However, as seen in Example 5.7, this will not happen very often.
@@
By Theorem 3.8, it suffices to show that the canonical group morphism $\operatorname{Pic}(j):\operatorname{Pic}(A)\rightarrow\operatorname{Pic}(B)$ given by $L\mapsto L\otimes_{A}B$ is injective where $j:A\rightarrow B$ is the inclusion map.
More generally, let $M$ and $N$ be arbitrary $A$-modules such that $M\otimes_{A}B\simeq N\otimes_{A}B$ as $B$-modules. By hypothesis, there exists a ring map $f:B\rightarrow A$ such that $f(a)=a$ for all $a\in A$. Then $f$ induces an isomorphism of $A$-algebras $B/I\rightarrow A$ which is given by $b+I\mapsto f(b)$ where $I=\operatorname{Ker}(f)$. In particular, the $A$-modules AA and $B/I$ are isomorphic. Then as $A$-modules we have $M\simeq M\otimes_{A}A\simeq M\otimes_{A}B/I\simeq(M\otimes_{A}B)\otimes_{B}B/I\simeq(N\otimes_{A}B)\otimes_{B}B/I\simeq N\otimes_{A}B/I\simeq N\otimes_{A}A\simeq N$. It is important to note that in the fourth isomorphism above, the assumption that $M\otimes_{A}B$ and $N\otimes_{A}B$ are isomorphic as “$B$-modules” is crucial. This completes the proof. ∎
@@
Remark 5.11.
Under the hypothesis of the above lemma, for any $A$-module $M$, then $M\otimes_{A}B$ has two $B$-module structures that are not identical in general: $b.(x\otimes b^{\prime})=x\otimes bb^{\prime}$ and $b\ast(x\otimes b^{\prime})=f(b)x\otimes b^{\prime}=x\otimes f(b)b^{\prime}$. But note that the functorial property of the Picard group allows us to consider only the first $B$-module structure here.
@@
The ring extension $A\subseteq A\oplus M$ has a retraction. In fact, the ring map $A\oplus M\rightarrow A$ given by $(r,m)\mapsto r$ is its retraction. Then apply Theorem 5.10. ∎
@@
But the ring extension $B\subseteq B\otimes_{A}B$ has a retraction. Then by Theorem 5.10, $\operatorname{\mathfrak{C}}(B,B\otimes_{A}B)=0$. ∎
@@
To prove the assertion, by Theorem 5.10, it suffices to show that the ring extension $R\subseteq RB$ has a retraction. First note that, since $A\setminus Z(A)\subseteq B\setminus Z(B)$, we have the extension of rings $T(A)\subseteq T(B)$. Hence, we really do have the ring $RB$.
@@
Let $A$ be a ring such that the extension $T(A)_{\mathrm{red}}\subseteq T(A_{\mathrm{red}})$ has a retraction, then using Theorem 5.10 and Corollary 5.6, we have the canonical isomorphism of groups $\mathfrak{C}(A)\simeq\mathfrak{C}(A_{\mathrm{red}})$. This can be viewed as a generalization of Corollary 5.8.
In the following example we will observe that Theorem 5.10 is in full generality. This means that it cannot be further generalized to split extensions. By a split extension we mean an extension of rings $A\subseteq B$ that splits as $A$-modules (i.e., there exists a morphism of $A$-modules $f:B\rightarrow A$ such that $f(a)=a$ for all $a\in A$). It is clear that split extensions are natural generalizations of ring extensions which have retractions. The following example also shows that there are extensions of rings $A\subseteq B$ such that $B$ is a free $A$-module and has a basis containing the unit element 1 (such extensions are special cases of split extensions), but $\operatorname{\mathfrak{C}}(A,B)\neq 0$.
@@
Acknowledgments. We would like to give sincere thanks to Professor Pierre Deligne who generously shared with us his very valuable and excellent ideas. The author would also like to thank Professor Melvin Hochster for helpful discussions.

### v3

@@ 3. Ideal class group of an extension vs Picard group
Definition 3.3.
Let $A\subseteq B$ be an extension of rings. The group $\mathscr{G}(A,B)$ modulo $H=\{Ax:x\in B^{\ast}\}$, the subgroup of principal invertible ideals of this extension, is called the ideal class group of this extension and is denoted by $\mathfrak{C}(A,B)$.
For any extension of rings $A\subseteq B$ we have the following short exact sequence of Abelian groups:
@@
The following result guarantees that the concept of the invertible ideal coincides with the classical concept of the invertible module:
@@
$\mathbf{(ii)}$ $L$ is a finitely generated projective $A$-module of rank 1 and $LB=B$.
@@
Note that Lemmas 3.4 and 3.6 also improve [8, Lemmas 2.2, 2.3].
@@
For any ring $A$ we have the following exact sequence of Abelian groups:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{C}(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Pic}(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Pic}(T(A)).}$
@@ 4. The geometry of line bundles
Let $L$ be an invertible module over a ring $A$. If $f:L\rightarrow A$ is a morphism of $A$-modules, then $f(x)y=f(y)x$ for all $x,y\in L$. In particular, the $A$-algebra $\bigoplus\limits_{n\in\mathbb{Z}}L^{\otimes n}$ is a commutative ring.
@@
The Picard group can be realized as the ideal class group:
@@
It follows from Theorems 3.9 and 4.2. As an alternative proof, in the proof of Theorem 4.2, we observed that for any ring AA then there exists an extension ring $A_{1}$ of $A$ such that the induced map $\operatorname{Pic}(A)\rightarrow\operatorname{Pic}(A_{1})$ is zero. Then by Theorem 3.9, we have $\operatorname{Pic}(A)\simeq\mathfrak{C}(A,A_{1})$. ∎
@@
We give a new proof to the following well known result:
@@
The Picard group does not necessarily commute with the “infinite” direct product of rings. However, we have the following result:
Theorem 4.6.
Let $(R_{i})$ be a family of rings. Then the group $\operatorname{Pic}(\prod\limits_{i}R_{i})$ can be canonically embedded in the group $\prod\limits_{i}\operatorname{Pic}(R_{i})$.
Proof.
We show that the natural morphism of groups $\operatorname{Pic}(\prod\limits_{i}R_{i})\rightarrow\prod\limits_{i}\operatorname{Pic}(R_{i})$ given by $L\mapsto(L\otimes_{R}R_{i})_{i}$ is injective. To see this, it suffices to show that every invertible $R$-module and more generally every finitely presented $R$-module $L$ is isomorphic to $\prod\limits_{i}L_{i}$ where $L_{i}=L\otimes_{R}R_{i}$ (recall that every finitely generated projective module is finitely presented). In fact, we show that the natural morphism of $R$-modules $\varphi:L\rightarrow\prod\limits_{i}L_{i}$ given by $x\mapsto(x\otimes 1)_{i}$ is an isomorphism. Indeed, first tensoring a finite presentation:
${R^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${L\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
of $L$ with $\otimes_{R}R_{i}$ to get a finite presentation for $L_{i}$:
${R_{i}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R_{i}^{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${L_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
Next taking direct products, the following exact sequence of $R$-modules is obtained:
${(\prod\limits_{i}R_{i}^{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\prod\limits_{i}R_{i}^{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\prod\limits_{i}L_{i})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
Then by using the fact that for any family of modules $(N_{k})$ over a ring $R$, the $R$-modules $(\prod\limits_{k}N_{k})^{\oplus n}$ and $\prod\limits_{k}N_{k}^{\oplus n}$ are canonically isomorphic for all natural numbers $n\geqslant 0$, we get the following commutative diagram of $R$-modules with exact rows:
[…]
So by the five lemma, the map $\varphi$ is an isomorphism of $R$-modules. ∎
Corollary 4.7.
Let $(R_{i})$ be a family of rings and $R=\prod\limits_{i}R_{i}$. Then $\operatorname{Pic}(R)=0$ if and only if $\operatorname{Pic}(R_{i})=0$ for all $i$.
Proof.
First note that for any idempotent $e$ of a commutative ring $R$, then $Re$ is a commutative ring with the identity element $e$, and we have a decomposition of rings $R\simeq Re\times R(1-e)$. Thus $\operatorname{Pic}(R)\simeq\operatorname{Pic}(Re)\times\operatorname{Pic}(R(1-e))$. Each factor ring $R_{k}$ is isomorphic to $R/(1-e)\simeq Re$ for some idempotent $e=(\delta_{i,k})_{i}\in R$ where $\delta_{i,k}$ is the Kronecker delta. Now if $\operatorname{Pic}(R)=0$, then $\operatorname{Pic}(R_{k})\simeq\operatorname{Pic}(Re)=0$ for all $k$. The reverse implication follows from Theorem 4.6. ∎
Corollary 4.8.
The Picard group of every direct product of local rings is trivial.
Proof.
Apply Corollaries 4.5 and 4.7. ∎
A weak version of Theorem 4.2 can be proved in an elementary fashion as follows:
Corollary 4.9.
Every ring has a ring extension whose Picard group is trivial.
Proof.
Every ring $R$ can be canonically embedded in $A:=\prod\limits_{M\in\operatorname{Max}(R)}R_{M}$. By Corollary 4.8, we have $\operatorname{Pic}(A)=0$. ∎
@@
By Theorem 4.2 or Corollary 4.9, there exists an extension ring $B$ of $A$ such that $\operatorname{Pic}(B)=0$. Then by Theorem 3.9, there exists an invertible ideal $L^{\prime}$ of the extension $A\subseteq B$ such that $L\simeq L^{\prime}$ as $A$-modules. Then by Lemma 4.13, $L=L_{k}$ for some $k$. ∎
The following result improves the related result in the literature:
Lemma 4.15.
Let $M$ be a finitely generated flat module and $N$ be a finitely generated projective module over a ring $R$ whose rank maps are the same, and let $f:M\rightarrow N$ be a surjective morphism of $R$-modules. Then $f$ is an isomorphism.
Proof.
Since $N$ is projective, we have a decomposition of $R$-modules $M\simeq N\oplus K$ where $K=\operatorname{Ker}(f)$. Since $M$ is a finitely generated flat $R$-module, $K$ is as well. Then, between the rank maps, we have the relation $\mathrm{r}_{M}=\mathrm{r}_{N}+\mathrm{r}_{K}$. Thus by hypothesis, $\mathrm{r}_{K}=0$. This shows that $K_{\mathfrak{p}}=0$ for all $\mathfrak{p}\in\operatorname{Spec}(R)$. Hence $K=0$. ∎
Lemma 4.16.
Let $e_{1},\ldots,e_{m}$ be idempotents of a ring $R$ such that $\operatorname{Spec}(R)=\bigcup\limits_{i=1}^{m}D(e_{i})$. Then there are orthogonal idempotents $g_{1},\ldots,g_{d}\in R$ ($d\leqslant 2^{m}-1$) such that for each $k$ there exists some $i$ with $g_{k}=g_{k}e_{i}$ and $\sum\limits_{k=1}^{d}g_{k}=1$.
Proof.
For each nonempty subset $S$ of $\{1,\ldots,m\}$, consider the idempotent $g_{S}:=(\prod\limits_{i\in S}e_{i})(\prod\limits_{j\notin S}(1-e_{j}))$. Clearly the idempotents $g_{S}$ are orthogonal and $g_{S}=g_{S}e_{i}$ for $i\in S$. We claim that $\sum\limits_{S}g_{S}=1$ where $S$ ranges in the set of nonempty subsets of $\{1,\ldots,m\}$. Since the idempotents $g_{S}$ are orthogonal, the element $\sum\limits_{S}g_{S}$ is idempotent. Then to prove $\sum\limits_{S}g_{S}=1$, it suffices to show that $D(\sum\limits_{S}g_{S})=D(1)$. For any idempotents $e,e^{\prime}\in R$ we have $D(e)\cup D(e^{\prime})=D(e+e^{\prime}-ee^{\prime})$. Using this and the orthogonality of the $g_{k}$, we get that $D(\sum\limits_{S}g_{S})=\bigcup\limits_{S}D(g_{S})$. Suppose there is some $\mathfrak{p}\in D(1)=\operatorname{Spec}(R)$ such that $\mathfrak{p}\notin\bigcup\limits_{S}D(g_{S})$. Then for each nonempty subset $S$ of $\{1,\ldots,m\}$ we have $g_{S}\in\mathfrak{p}$. But $e_{k}\notin\mathfrak{p}$ for some $k$. For $S=\{k\}$ we have $g_{S}=e_{k}(\prod\limits_{j\neq k}(1-e_{j}))\in\mathfrak{p}$. So there exists some index $\ell$ with $\ell\neq k$ such that $1-e_{\ell}\in\mathfrak{p}$. Then take $S=\{k,\ell\}$, we have $g_{S}=e_{k}e_{\ell}(\prod\limits_{j\neq k,\ell}(1-e_{j}))\in\mathfrak{p}$ and so $1-e_{j}\in\mathfrak{p}$ for some $j\neq k,\ell$. Continuing this process, we conclude that for $S=\{1,\ldots,m\}$, we have $g_{S}=\prod\limits_{i=1}^{m}e_{i}\notin\mathfrak{p}$ which is a contradiction. ∎
Theorem 4.17.
The Picard group of every zero-dimensional ring is trivial.
Proof.
Let $R$ be a zero-dimensional ring. By Theorem 2.1, we have a canonical isomorphism of groups $\operatorname{Pic}(R)\simeq\operatorname{Pic}(R_{\mathrm{red}})$. We know that $R_{\mathrm{red}}$ is a von Neumann regular (absolutely flat) ring. Hence, to prove the assertion, it suffices to show that the Picard group of every von Neumann regular ring $R$ is trivial. Let $L$ be an invertible $R$-module. So the canonical morphism of $R$-modules $\widehat{L}\otimes_{R}L\rightarrow R$ is an isomorphism. Then we may write $1=\sum\limits_{i=1}^{m}f_{i}(x_{i})$ where $f_{i}\in\widehat{L}$ and $x_{i}\in L$ for all $i$. Since $R$ is von Neumann, for each $a\in R$ there exists some $b\in R$ such that $a=a^{2}b$. Then $e=ab$ is idempotent and $Ra=Re$. Therefore, for each $i$, there exists an idempotent $e_{i}\in R$ such that $f_{i}(x_{i})R=e_{i}R$ and $\operatorname{Spec}(R)=\bigcup\limits_{i=1}^{m}D(e_{i})$. By Lemma 4.16, there are orthogonal idempotents $g_{1},\ldots,g_{d}\in R$ such that for each $k$ there exists some $i$ for which $g_{k}=g_{k}e_{i}$ and $\sum\limits_{k=1}^{d}g_{k}=1$. Clearly the map $h:g_{k}L\rightarrow g_{k}R$ given by $g_{k}x\mapsto g_{k}f_{i}(x)$ is a morphism of $g_{k}R$-modules (note that $g_{k}R$ is a ring with the identity $g_{k}$). This map is surjective, because $g_{k}r=g_{k}e_{i}r=g_{k}f_{i}(x_{i})r^{\prime}=g_{k}f_{i}(r^{\prime}x_{i})$. Clearly $g_{k}L\simeq L/(1-g_{k})L\simeq L\otimes_{R}R/(1-g_{k})$ is an invertible module over $g_{k}R\simeq R/(1-g_{k})$. Then by Lemma 4.15, $h$ is an isomorphism of $g_{k}R$-modules (and so it is an isomorphism of $R$-modules). Since the $g_{k}$ are orthogonal with summing 1, we have the isomorphisms of $R$-modules $L=\sum\limits_{k=1}^{d}g_{k}L\simeq\bigoplus\limits_{k=1}^{d}g_{k}L\simeq\bigoplus\limits_{k=1}^{d}g_{k}R\simeq\sum\limits_{k=1}^{d}g_{k}R=R$. Hence, $\operatorname{Pic}(R)=0$. ∎
In particular, Picard groups of finite rings, Artinian rings, and Boolean rings are trivial.
@@ 5. The ideal class group behavior with a tower of extensions
where the rows are exact (see Theorem 3.9) and the second and third vertical maps are the canonical isomorphisms (see Theorem 2.1). The above diagram is also commutative, since the commutativity of the left-hand side follows from Corollary 3.7, and the right-hand side is clear. Then by the five lemma, the first vertical map is an isomorphism. ∎
@@
where the rows are exact and the second vertical map is an isomorphism (see Theorem 2.1). Hence, the canonical map $\operatorname{\mathfrak{C}}(A)\rightarrow\operatorname{\mathfrak{C}}(A_{\mathrm{red}})$ is injective. However, unlike the Picard group, the canonical injection $\operatorname{\mathfrak{C}}(A)\rightarrow\operatorname{\mathfrak{C}}(A_{\mathrm{red}})$ is not necessarily an isomorphism in general (see Example 5.7).
@@
where the rows are exact (see Corollary 3.11) and the second vertical map is an isomorphism (Theorem 2.1). If the third vertical map is injective, then by the five lemma, the canonical group map $\operatorname{\mathfrak{C}}(A)\rightarrow\operatorname{\mathfrak{C}}(A_{\mathrm{red}})$ will be an isomorphism. However, as seen in Example 5.7, this will not happen very often.
@@
By Theorem 3.9, it suffices to show that the canonical group morphism $\operatorname{Pic}(j):\operatorname{Pic}(A)\rightarrow\operatorname{Pic}(B)$ given by $L\mapsto L\otimes_{A}B$ is injective where $j:A\rightarrow B$ is the inclusion map. By hypothesis, there exists a ring map $f:B\rightarrow A$ such that $f\circ j:A\rightarrow A$ is the identity map. Then by the functorial property of the Picard group, we have $\operatorname{Pic}(f\circ j)=\operatorname{Pic}(f)\circ\operatorname{Pic}(j):\operatorname{Pic}(A)\rightarrow\operatorname{Pic}(A)$ is the identity map. This shows that $\operatorname{Pic}(j)$ is injective. ∎
@@
The ring extension $A\subseteq A\oplus M$ has a retraction. In fact, the ring map $A\oplus M\rightarrow A$ given by $(r,m)\mapsto r$ is its retraction. Then apply Corollary 5.10. ∎
@@
But the ring extension $B\subseteq B\otimes_{A}B$ has a retraction. Then by Corollary 5.10, $\operatorname{\mathfrak{C}}(B,B\otimes_{A}B)=0$. ∎
@@
To prove the assertion, by Corollary 5.10, it suffices to show that the ring extension $R\subseteq RB$ has a retraction. First note that, since $A\setminus Z(A)\subseteq B\setminus Z(B)$, we have the extension of rings $T(A)\subseteq T(B)$. Hence, we really do have the ring $RB$.
@@
Let $A$ be a ring such that the extension $T(A)_{\mathrm{red}}\subseteq T(A_{\mathrm{red}})$ has a retraction, then using Corollaries 5.10 and 5.6, we have the canonical isomorphism of groups $\mathfrak{C}(A)\simeq\mathfrak{C}(A_{\mathrm{red}})$. This can be viewed as a generalization of Corollary 5.8.
In the following example we will observe that Corollary 5.10 is in its full generality. This means that it cannot be further generalized to split extensions. By a split extension we mean an extension of rings $A\subseteq B$ that splits as $A$-modules (i.e., there exists a morphism of $A$-modules $f:B\rightarrow A$ such that $f(a)=a$ for all $a\in A$). It is clear that split extensions are natural generalizations of ring extensions which have retractions. The following example also shows that there are extensions of rings $A\subseteq B$ such that $B$ is a free $A$-module and has a basis containing the unit element 1 (such extensions are special cases of split extensions), but $\operatorname{\mathfrak{C}}(A,B)\neq 0$.
@@
Acknowledgments. We would like to give sincere thanks to Professors Brian Conrad, Pierre Deligne, and Melvin Hochster who generously shared with us their very valuable and excellent ideas.